OpenStudy (john_gardner):

WILL GIVE 4 MEDALS AND FAN IF YOU ANSWER PLEASE! To simplify the radical below, which of the following expressions would be multiplied by the radical? 2√15/(-4-√12) A.(-4+√12)/(-4+√12) B. (-4+√12) C. 1/(-4+√12) D. (-4+√12)/(-4-√12)

OpenStudy (john_gardner):

@agent0smith @debpriya @Ethan01 helpppp lol

OpenStudy (john_gardner):

@Zyberg @Dragontears1128 @PartyShotz2409 help?

OpenStudy (john_gardner):

someone*

OpenStudy (dragontears1128):

https://www.wolframalpha.com here is a website that can help you as well if your stuck on math

OpenStudy (dragontears1128):

hope that helps

OpenStudy (dragontears1128):

i us it and https://mathway.com/

OpenStudy (john_gardner):

I QUIT oml im just gonna get it wrong

OpenStudy (mathmate):

To rationalize a denominator of the form \((a+\sqrt b)\) we multiply by the conjugate, \((a-\sqrt b)/(a-\sqrt b)\) Example, to rationalize \(\frac{5+\sqrt{5}}{3-\sqrt{7}}\) we multiply by: \(\frac{3+\sqrt{7}}{3+\sqrt{7}}\) to get \(\frac{5+\sqrt{5}}{3-\sqrt{7}} \times \frac{3+\sqrt{7}}{3+\sqrt{7}}\) \(=\frac{(5+\sqrt{5})\times(3+\sqrt{7})}{(3-\sqrt{7})\times (3+\sqrt{7})} \) \(=\frac{(5+\sqrt{5})\times(3+\sqrt{7})}{3^2-\sqrt{7^2}} \) \(=\frac{(15+3\sqrt{5}+5\sqrt{7}+\sqrt{35})}{3^2-\sqrt{7^2}} \) =\(\frac{(15+3\sqrt{5}+5\sqrt{7}+\sqrt{35})}{2}\)